OpenStudy (anonymous):

A simple pendulum on Earth has a period of 6.0 seconds. What would the period of this pendulum be on Jupiter, if the acceleration of gravity on the surface of Jupiter is 26 m/s ? A. 2.3 sec B. 3.7 sec C. 6.0 sec D. 15.6 sec

OpenStudy (anonymous):

The restoring force of a pendulum \[-mg \sin \theta = ma\] For small angles, \[\sin \theta = \theta\] so \[a=\frac{\mathrm{d}^2v}{\mathrm{d}t^2}=\ell \frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}=-g\theta\] giving the simple diff eq \[\frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}+\frac{g}{\ell}\theta=0\] with the solution \[\theta=\theta_{max} \sin(\omega t)\] \[ \text{angular frequency} \ = \omega = \sqrt{\frac{g}{\ell}}\] \[ \text{Period}=T=\frac{2\pi}{\omega} = 2\pi \sqrt{\frac{\ell}{g}}\] Solve for l - the length of the pendulum - by plugging in the period on Earth along with earth's gravity, then use the found length and the gravity of Jupiter to find the period on Jupiter ^_^

OpenStudy (anonymous):

The period should be shorter on Jupiter.

OpenStudy (anonymous):

3.7 sec? @AllTehMaffs

OpenStudy (anonymous):

yup yup ^_^

OpenStudy (anonymous):

@Partycool nice!